OpenStudy (anonymous):

(someone please helpppp! D: it's the last day ) 2. The table shows the relationship between variables. Which selection describes the relationship? x y 1 1 2 -3 3 -7 4 -11 a. increasing; linear b. increasing; non-linear c. decreasing; linear d. decreasing; non-linear

OpenStudy (anonymous):

please help D: I really need it

OpenStudy (misty1212):

hi! as x goes up does y go up or down?

OpenStudy (anonymous):

when you increase x y decreases 4 in each step so it is decreasing linearly

OpenStudy (misty1212):

or just go with the answer given above...

OpenStudy (bohotness):

which one do you think is your answer?

OpenStudy (anonymous):

o_oohh wait what?

OpenStudy (bohotness):

what one do yuo think is the answer?

OpenStudy (anonymous):

ummm C?

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

*sigh*

OpenStudy (bohotness):

yes

OpenStudy (bohotness):

:d

OpenStudy (anonymous):

yay! can you help me on a few more please? I really need to get this done D:

OpenStudy (bohotness):

:DDDD

OpenStudy (bohotness):

sure

OpenStudy (anonymous):

The ordered pairs (1,3), (2,9), (3,27), (4,81), and (5,243) represent a function. what is a rule that represents this function? a. y+x^3 b. y=3x c. y+ 3^x d. y= x+3

OpenStudy (bohotness):

which one do you think is the answer?

OpenStudy (anonymous):

idk!!

OpenStudy (anonymous):

like umm *guesses* c?!?!

OpenStudy (bohotness):

i think you right

OpenStudy (anonymous):

am I? :o

OpenStudy (anonymous):

are you sure? :o because I need to pass all of this :(

OpenStudy (bohotness):

wait what school are you doing this from

OpenStudy (anonymous):

connexus (Connections Academy )

OpenStudy (bohotness):

O..O

OpenStudy (anonymous):

I am at my computer desk freaking out and stressing!

OpenStudy (bohotness):

I am in k12 online school

OpenStudy (anonymous):

oooooooo :o

OpenStudy (bohotness):

go to www.brainly.com

OpenStudy (bohotness):

:) they are like us but a lot differnet

OpenStudy (anonymous):

does it give me answers?

OpenStudy (anonymous):

@Colere252346

OpenStudy (anonymous):

@misty1212

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

no it didn't yet :/ @bohotness

OpenStudy (bohotness):

?????

OpenStudy (anonymous):

no one has seen my question and I'm stressing over this !

OpenStudy (bohotness):

@AnswerMyQuestions

OpenStudy (anonymous):

who that? :O are they on D:

OpenStudy (anonymous):

@aleroth

OpenStudy (anonymous):

ughh! no one is gonna help :(

OpenStudy (anonymous):

The ordered pairs (1,3), (2,9), (3,27), (4,81), and (5,243) represent a function. what is a rule that represents this function? a. y+x^3 b. y=3x c. y+ 3^x d. y= x+3

OpenStudy (anonymous):

it's C

OpenStudy (anonymous):

you were right >_> @bohotness

OpenStudy (anonymous):

Imma make a new question

OpenStudy (bohotness):

okay i help

OpenStudy (bohotness):

:d

OpenStudy (bohotness):

lol

OpenStudy (bohotness):

k